I really could use some help with this problem.
(SUM 14)(-0.4i+4)
i=1
I would appreciate any help that I can get.

Answer by oscargut(891)

Sum (from i=1 to i=14)of (-0.4i+4)
there are 2 ways of doing it
1) writing all terms
Sum = S = [-0.4(1)+4]+[-0.4(2)+4]+.......[-0.4(14)+4]=14
2) using that sum(from i=0 to i=n)of (i)=i(i+1)/2
(Example sum (from i=1 to i=5) of (i) is 5(6)/2=15)

then the sum S= Sum (from i=1 to i=14)of (-0.4i+4)=(using properties of a finite sum)=-0.4[sum (from i=1 to i=14) of (i)]+14(4)=
-0.4[14*15/2]+56=-0.4(105)+56=-42+56=14
Answer: S=14